A baseball player is running from the second base to the third base at 20ft/sec. At what rate is his

distance from the home plate changing when he is 30ft from the third base. The baseball diamond

is a square 90ft on aside.


1
Expert's answer
2022-04-17T15:13:22-0400

The picture represents a baseball diamond and a player.



The player is located at the point "E". He is between the third "(C)" and the second base ("B" ). The arrow shows the velocity vector. "|CE|=30ft". "|AE|" is the distance between the home plate and the player. Using the Pythagorean theorem, we get: "|AE|^2=|AB|^2+|EB|^2=90^2+60^2={11700}". "|AF|" denotes the distance between the home plate and the player, when he will pass a small distance "|EF|". "|EF|=v\Delta t", where "\Delta t" is a small period of time. "|AF|^2=|AB|^2+(|FE|+|BE|)^2=90^2+(v\Delta t+60)^2". It is enough to find the derivative of "|AF|-|AE|" with respect to "\Delta t" and substitute "\Delta t=0." It will show how fast does the distance change. "|AF|'-|AE|'=|AF|'=\frac{1}{|AF|}v(v\Delta t+60)". After setting "\Delta t =0" we get: "60v\frac{1}{|AE|}=\frac{1}{\sqrt{11700}}60\cdot20\approx11.09". Thus, the distance between the home plate and the player changes with the velocity "11.09ft\/sec". This is the velocity, when the player is at the point "E" . It changes with time.